# TabSketchFM: Sketch-based Tabular Representation Learning for Data Discovery over Data Lakes

### v1

@@ 3. TabSketchFM
1. Numerical sketches. For each column, we record the number of NaNs and the number of unique values (normalized by the number of rows in the table). For string columns, we also record cell width in bytes, since that can be an important determinant in governing if the column is likely to be a join column. For instance, very long strings are unlikely to be join candidates. When possible, we convert date columns into timestamps and then treat them as numeric columns. For numeric columns, we additionally compute a percentile sketch, and add the mean, the standard deviation, and minimum and maximum values. All these features are encoded as individual elements of a single input vector we call the numerical sketch.
@@
2. MinHash sketches. For each column, we compute a MinHash signature of cell values. [^2] In our implementation, we use datasketch library (Zhu and Markovtsev, 2017) to compute MinHash. As shown in Fig. 1, the cell value of Austria Vienna is used as a single element in the set passed to the MinHash computation. [^3] While it does not make sense to convert float or integer columns into MinHash signatures, we do so because it is often difficult to tell if a column is truly a float or an integer, or it really is a categorical value. For string columns, we also compute another MinHash signature for the tokens within the column. The rationale for this second MinHash is that the tokens can sometimes capture its semantics (e.g., if a token street appears in two different columns, they maybe both about address, which is useful if one considers a union of two tables that do not necessarily share cell values). As shown in Fig. 1, the cell values of Austria and Salzburg get tokenized and used as two different elements of the set passed to the MinHash computation. We concatenate both MinHash sketches into a single input vector for string columns. For numerical and date columns, only the MinHash for the cell values is included in the input vector.
3. Content snapshot. Row-based information could be crucial in understanding the table semantics for data discovery (Khatiwada et al., 2023a). We concatenate values within each row to form a string and create MinHash signature for each of them. For instance, we concatenate the last row of table in Fig. 1, surrounded by a green rectangle, into "Quarterly Austria Salzburg 800,000 28/03/23" and hash it as another minhash style sketch.
Now that we have these sketches, which are arrays of numbers, a key question is how to adapt BERT’s architecture to consume these numeric values. Transformer models take sentences as input, and internally break it into different inputs such as the tokens of a sentence [^4] The BERT tokenizer sometimes breaks a single word into multiple tokens such that unknown words in its vocabulary are handled correctly., the position of tokens in the sentence, and a scalar type indicating if the token belongs to the first sentence or the second. We leverage this mechanism to pass the table-specific scalar inputs as shown in  Fig. 1 to the BERT Embedding Layers. For vector inputs (e.g., numerical sketch, MinHash sketch), we expand the notion of BERT Embeddings to add some linear layers, as described below.
TabSketchFM’s specialized layers for processing tabular input are shown in the light gray box, middle-right of Fig. 1. The token Price for the Price column is used as an example for illustrative purposes. All other tokens belonging to the table description and other columns would undergo a similar process to create a single embedding per token.
1. Token Embedding Layer. The embedding layer of the BERT uncased model (Devlin et al., 2019) contains numerical embeddings for each word in its vocabulary based on extensive training with a large corpus, such that, for instance, the embedding of Price might be closely associated with that of House because they co-occur often in text. Our model’s token embedding layer is initialized with the weights of the BERT model’s embedding layer to leverage such information present in natural language, but the weights are allowed to be changed further while pretraining to include co-occurrence information in table-specific corpora. As shown in Fig. 1, the token Price would be mapped from its position in the vocabulary of the model (also pre-populated with BERT’s vocabulary) into its hidden state using weights from this token embedding layer.
@@
2. Token Position Embedding Layer. Each column is analogous to a sentence in a paragraph. Hence, we re-purpose token positional embeddings to reflect a token’s position within a column name. For instance, as shown by the light gray scalar input in Fig. 1, the token position of a token Area is 1 because it is the second token in the cell (after Reference).
@@
3. Column Position Embedding Layer. Columns themselves have positions and are encoded through a new embedding layer which can range from 1 to the total number of columns, [^5] Position 0 is reserved for any tokens in the table description. as shown by the white scalars in Fig. 1. The rationale for including column positions is that they sometimes do have semantics; e.g., a street address next to the city, followed by a zipcode. Of course table semantics do not change as a function of column order; nevertheless, we included position in case it helps the model understand column semantics, with the assumption that the attentional mechanism would disregard order if necessary.
@@
4. Column Type Embedding Layer. Column types of string, date, integer, or float are encoded through another embedding. In order to determine column type, we made a best-case effort to parse the first 10 values of each column as dates, integers, or floats and defaulted to string if we could not convert them. In mixed-type columns, this can yield poor results, but at the very worst, at least one of the types was assigned to these columns. Column types are shown in Fig. 1 in orange.
5. MinHash Sketch Linear Layer. MinHash sketches and the content snapshots were passed through a linear layer to encode them into the same hidden state dimensions of the existing transformer layers in Bert.
6. Numerical Sketch Linear Layer. Numerical sketches are similarly passed through a linear layer to encode them.
As shown in Fig. 1, for each column token, we combined the hidden states of its token embedding, token position embedding, column position embedding, column type embedding, MinHash sketch embedding, and numerical sketch embedding using summation. This summation is similar to the summation of token embeddings and position embeddings in the original BERT model. Content snapshot signatures are summed with table description tokens because the content snapshot describes the entire table’s content. For instance, for the tokens Residential, Property and Prices, the content snapshot is passed into the linear layers to create a hidden state that is combined with the other embeddings for those tokens. Once each token has been summed with its other sketches, positional and type encoding, it is passed to the rest of the encoder and attention layers from BERT. The attention layers are crucial for contextualizing the embeddings based on what other columns, values, and table descriptions exist in the table. As shown in the figure for Price, the attention mechanism learns $N^{2}$ attention weights between all $N$ tokens for a given table. This step will consider Price in relation to all other columns as well as table descriptions. We used the usual 12-layer BERT encoder model with self attention from HuggingFace [^6] https://huggingface.co/docs/transformers/model_doc/bert once we have unified the hidden states from the embedding layers and the linear layers.
4. Pretraining
@@
Dataset. Existing works generally use non-public data (Badaro et al., 2023) or web tables (Lehmberg et al., 2016) to pretrain the models. Unlike enterprise data, web tables often have few rows and columns to ease human consumption (Cafarella et al., 2008) and they focus on entities popular on the web (e.g., countries). Their performance may not generalize to enterprise data, where the tables have a large number of rows, the entities are often domain-specific, contain cryptic code words, and have a lot of numerical information. Therefore, we created a de-duplicated pretraining dataset of $197,254$ enterprise-like tables (CSVs) from CKAN and Socrata that are legally cleared to have the open licenses. This pretrain dataset contains $2234.5$ rows and $35.8$ columns in each table on average. About $66\%$ of columns were non-string, resembling an enterprise datalake.
@@
Data Augmentation. In the image processing literature (Acharya et al., 2006), different transformations are applied over an original image to generate new image samples for training. This makes the trained model robust on different variants of the same image. We want the model to be robust to the order of columns such that shuffling columns in a table does not impact its semantics. So, we created three different versions of the table, by changing the column order, which in turn, changed the table’s content snapshot (see Section 3). This increases the total number of pretraining tables to 290,948 and for each of them, we create signatures as explained in Section 3.
@@
Figure 2. A combination of whole column masking, and MLM probability masking to generate up to 5 examples. The box on the left hand side is the table description and the one on the right hand side is the column names
In our scenario, for each table, we mask a single column name such that all tokens corresponding to that column are masked. This is analogous to the natural language literature’s whole word masking where all tokens corresponding to a word are masked. In addition, we used MLM probability to mask the tokens in the table description as well. As shown in Fig. 2, we generate several samples from a single table. Note that each masking gives us an example. For small tables with less than $5$ columns, we mask each of its columns one after another. However, if a table has a large number of columns, this produces a lot of examples for the same table, over-representing it. So, for large tables with more than $5$ columns, we randomly select five columns and mask them. Following this strategy over the augmented pretraining dataset, we get $730,553$ examples in training, $54,430$ examples in validation, and $58,141$ examples in test sets. We provide the experimental setup for pretraining in Section 6.
5. Finetuning Datasets
Table 1. Cardinality of all the datasets in LakeBench.
@@ 5. Finetuning Datasets
After pretraining, we need finetuning to adapt the pretrained model to particular tasks of identifying relevant tables such as unionability, joinability, or subset. Existing finetuning datasets target other tasks such as table metadata prediction (Deng et al., 2020), question answering over tables (Yin et al., 2020), and so on, and none focus on relevant table understanding. Therefore, we create LakeBench, a collection of finetuning datasets to identify relavant table pairs using multiple data sources such as open government data from CKAN and Socrata, economic data from the European Central Bank (Eurosystem Data, 2023), Spider (Yu et al., 2018), and synthesized data from large knowledge graphs such as Wikidata (Vrandečić and Krötzsch, 2014). Crucially, we ensure that if a table is a part of the pretraining dataset, we never use it as a part of any finetuning dataset.
LakeBench covers Unionability (Section 5.1), Joinability (Section 5.2) and Subset (Section 5.3) tasks, and a wide range of problem types: binary classification, regression, and multi-label classification. Each benchmark comes with train, test, and validation sets where each example contains a table pair and its label. Table 1 summarizes the tasks and dataset cardinalities.
5.1. Unionability
Two tables A and B are unionable if a subset of their columns are unionable (Nargesian et al., 2018), and fully unionable if all their columns are unionable. A column $c_{1}$ from Table A is considered unionable with a column $c_{2}$ from Table B if they contain values that are drawn from the same domain. While overlap between values in the columns can indicate that they belong to the same domain, columns with few or no overlapping values may also map to the same semantic type or concept. For instance, two columns titled “movie" and “film" might have different movie names in them with no overlap, but they can still be unionable because they map to the same semantic type. Generally, we take a union of a new table with an existing table to expand it vertically (i.e., add more rows) (Khatiwada et al., 2023a). LakeBench contains three datasets for the unionability task: one adapted from existing table search benchmarks (Khatiwada et al., 2023a; Nargesian et al., 2018), one synthesized from Wikidata, and one from ECB data.
1. TUS-SANTOS
Nargesian et al. (2018) created the TUS Benchmark to evaluate the performance of table union search methods over an open data lake. They use $32$ large seed tables from Canada’s Open Government Data Portal (of Canada, 2023) and U.S. Government’s Open Data (Government, 2023). The seed tables are selected from unique domains such that they are not unionable with each other. Then, they generate $5000$ smaller tables by randomly sampling the rows and columns from each seed. The smaller tables generated from the same seed are unionable to each other and the tables from different seeds are not unionable. Khatiwada et al. (2023a) created SANTOS Benchmark by adapting the same benchmark creation technique (Nargesian et al., 2018), but preserving the context and the key entity column during the split. For example, if a seed table contains columns (“person", “birthplace", and “birth year"), the TUS approach might result in a table with (“birthplace" and “birth year"), while the SANTOS approach ensures that the “person" column is preserved (Khatiwada et al., 2023a). These benchmarks were constructed for a search task where the requirement is to retrieve unionable tables from a data lake given a query table. To curate a benchmark for finetuning the pretrained model for the table unionability task, we sample unionable and non-unionable table pairs from TUS and SANTOS and create a binary classification dataset.
2. Wiki Union
To curate a larger dataset along with ground truth unionability assessment labels, we use the process depicted in Fig. 3. We take a structured knowledge base (KB) with a SPARQL (Sparql data, 2023) endpoint, and create a collection of tables describing entities in the input knowledge base. The process starts with creating a “profile" of the ontology of the KB. The profile consists of a list of all the classes, the properties of each class, and statistics about the number of instances of each class, the number of values for each property, as well as the data types. This profile is then used to specify a configuration that determines the characteristics of the resulting tabular data lake. The characteristics include: domain (i.e., types of entities), inclusion and prevalence of different data types (e.g., numerical or categorical), the minimum and maximum number of rows and columns in tables, the prevalence of tables that have the same schema, the prevalence of ambiguous entity labels, the amount of noise introduced (if any), the prevalence of tables with null values, and the maximum number of tables about the same entity type. The specification is then used to generate a raw collection of tables. The result is a possibly very large collection of tables, along with ground truth mappings to the KB.
Figure 3. Using a knowledge base to curate a tabular data collection for table union and join benchmarks
Figure 4. Sample subsets of four tables from our Wikidata-based table union and join benchmark (top), along with ground truth labels and mappings (bottom)
We use Wikidata (Vrandečić and Krötzsch, 2014) as the source KB and curate a tabular data collection. The resulting collection consists of large tables with mostly numerical columns, to resemble tabular data found in enterprise data lakes. Figure 4 shows example tables from this collection. The first column in each table contains the labels of entities in the KB. Every other column represents either a relation or a property of the entity. Relation columns contain labels of entities while property columns contain literal values. Some tables may contain null values in some of their rows and one of their columns. The configuration consisted of a subset of Wikidata classes along with only their numerical properties, in addition to a random subset of classes along with all their relations and properties. The configuration specified between 2 and 8 columns, and between 20 and 120 rows. As the classes in Wikidata are quite large and may result in a large number of tables and an unbalanced collection, we limit the number of rows in all the tables for each class to 3,000, and limit the number of tables with the same schema to 20. The raw collection consists of $46,521$ tables, with $3,157,781$ mappings of cell values to $1,317,724$ unique entities, $72,458$ property mappings, and $53,087$ column to concept mappings.
To curate the Wiki Union dataset, we designate two tables as fully unionable if they are about the same concept and all their columns map to the same properties in the KB as per the ground truth mappings. The first two tables in Fig. 4 are such fully unionable tables. We create two kinds of negative labels: a) tables in which columns map to the same properties but are about different entities; and b) tables with the same number of columns but not all of their columns map to the same properties. Tables QCXMIM62QXN0.csv and GSW089XWBCBS.csv in Fig. 4 are negative examples, with both tables having columns that map to area (P2046) and population (P1082), but the tables are about different types of entities (as shown in table cta_gt.csv) and not unionable.
3. ECB Union
The European Central Bank (ECB) distributes detailed information about the economy across the European Union (Eurosystem Data, 2023), organized into $74$ datasets, each dedicated to a different topic. Each dataset contains a set of time series tables, with varying features from a given dimension. Figure 5(a) shows an example from two datasets that are related: residential property prices (RESR) with over 500 tables, and residential property valuation (RPV) with 232 tables. Each table specifies the values for each dimension in the dataset, for instance, AT stands for Austria, AT130 stands for Vienna, EC stands for everything except Vienna, etc. For the ECB union benchmark, we created a regression task by pairing tables such as the table for Vienna, new flats as shown in the figure with the table for outside Vienna, new flats and the table for outside Vienna, all flats, with the assumption that the plausibility of the union depends on the number of changed values across the dimensions. Therefore, we construct ECB Union benchmark by ranking each pair by how many dimensions differ, which varies from 1 to 12 in our benchmark dataset. This desired ranking forms a regression task.
Figure 5. (a) Example Datasets from ECB (b) Spider-OpenData Join benchmark data generation: Once a valid join column is selected, the table is sorted based on the join column, and then divided into four quadrants to identify positive (horizontal) and negative (diagonal) examples.
5.2. Joinability
We consider a column from Table A as joinable with a column from Table B if the two columns map to the same semantic type and have overlapping values. We join a new table to an existing table to expand it horizontally (i.e. to add more features or columns) (Zhu et al., 2019). LakeBench contains four datasets for the joinability task; two created from Wikidata, another which was designed from Spider and CKAN/Socrata; and the last reflects existing joins from ECB.
1. & 2. Wiki Jaccard and Wiki Containment
We use the same corpus of Wikidata tables we used for Wiki Union to derive a benchmark for table joins. For this benchmark, we use the cell-entity (CE) mappings in the ground truth mappings and assign joinability scores to pairs of columns in the collection. The score is either the Jaccard similarity (size of intersection over the size of the union) across sets of CE mappings, or the minimum containment ratio across the sets of CE mappings which indicates an overlap in the entities in those columns and so a potential for joining. Tables T1GB2LNKTOXP.csv and SW4T7EFEI2DK.csv in Fig. 4 are examples of such tables in our benchmark, because they share a number of values in their first column that map to the same knowledge base entity. These two tasks are modeled as regression tasks.
3. Spider-OpenData Join
To create this benchmark, we used two data sources, Spider (Yu et al., 2018) and CKAN/Socrata open government data. Spider is a large-scale human-annotated text-to-SQL dataset annotated by 11 Yale students. It comes with 10K questions and 200 databases with multiple tables covering 138 different domains. Within each database, joinability is clearly identified via primary/foreign key relationships. Due to the relatively smaller number of tables per database, we were able to generate only a small number of join examples.
To ensure that we have enough samples for training and testing various models, we also used CKAN (CKAN data, 2023) and Socrata (Socrata data, 2023) open government data. However, this dataset did not come with manually annotated joins, so in the spirit of TUS (Nargesian et al., 2018) and SANTOS (Khatiwada et al., 2023a), we synthesized joins. Figure 5(b) illustrates this benchmark creation. For every table with at least 2 columns, 1) we select a join column at random from the set of columns with mostly unique values that are not floats, 2) we sort the table based on the join column, and 3) we divide the table around the join column into four quadrants. The top two quadrants and the bottom two quadrants are good candidates for positive joinable tables. Adjacent quadrants share the same join column and hence it is considered a true positive join. To create negative examples, we pair quadrants on the diagonal to form two pairs of examples, (top left, bottom right) and (bottom left, top right). We also ensure that these negative examples are true negatives and they do not share any values across the join column.
Figure 6. Pair creation for $S_{1}$ from CKAN/Socrata. To ensure positive and negative subset pairs differ by the same amount in the number of rows, $S_{1}$ is paired with the union of $S_{1}$, $S_{2}$, $S_{3}$ for a positive pair and $S_{2}$, $S_{3}$, $S_{4}$ for a negative pair.
4. ECB Join
As shown in Figure 5(a), 74 datasets in ECB have several shared dimensions (a total of 56 dimensions in all, with some tables sharing as many as 18 dimensions). Many of those datasets such as RESR (residential property prices) and RPV (residential property valuations) are related on multiple dimensions. For the join task, we collapsed across all tables within a given dataset into a single table, just so we could evaluate joins efficiently. Many tables are therefore large; the largest one has over 20 million rows. For each pair of tables, we computed joins on all shared dimensions to see if the result returned any rows. If it did, we recorded the dimensions on which the join was possible to model it as a multi-label classification problem. If the tables shared dimensions but a join resulted in no rows, this was recorded as another label (i.e., no joins are possible). The total dataset had 1780 pairs for training, 222 for valid and 223 for test. This multi-attribute join benchmark is modeled as a multiclass classification task. Although this is a relatively small dataset for finetuning, we include it as it specifies multi-attribute joins on very large, realistic tables, which is difficult to construct synthetically. It is also a measure of how much data is actually needed to finetune a model.
Figure 7. Example Table Subsets from CKAN Subset
5.3. Table Subsets
Table A is a subset of Table B if the rows of Table A are contained in B. The subset task is a useful one for governance as it is often useful to trace the provenance of data across tables in a data lake, and to find potential copies of a table (with some modifications). Table subset is defined as a binary classification task. We provided the models with the column names in the table, which meant that positive and negative examples had the exact same schema, but differed in values. In LakeBench, we provide a subset benchmark created using tables from CKAN/Socrata.
CKAN Subset
The subset problem becomes challenging when the schemas of the positive and negative examples are exactly the same. But random pairs of tables in the CKAN/Socrata data are most likely to have different schemas. So, we adopted a strategy shown in  Fig. 6. Each table greater than 100 rows was partitioned into 4 equal subsets, $S_{1-4}$. Each subset $S_{i}$ was paired with a table composed of $S_{i}$ and with two other subsets drawn randomly (e.g., $S_{2}$,$S_{3}$ in Fig. 6) for a positive example (shown in green), and paired with a table composed of all other subsets (i.e., $S_{k\neq i}$, shown in red) for a negative example. With this strategy, the number of rows differential for positive and negative pairs could not be used as a signal for whether a pair of tables was a subset or not. Much of the CKAN/Socrata data is de-normalized (Usta and Salihoglu, 2023), so it contains many repeating values in certain columns (e.g. name of a department). This means that negative examples often overlap in some of the columns, which makes the subset problem more difficult. Figure 7 shows example table subsets from CKAN Subset.
6. Experiments
Now we empirically address the following research questions:

### v3

@@ III-A Sketches Preparation
Numerical Sketches. For each column, we extract a set of statistical features that provide insight into its characteristics. These features include the number of unique values and of nan values, each normalized by the total number of rows. For string columns, we record the average cell width in bytes, as this can impact the likelihood of the column being used as a join key. Notably, very long strings are unlikely to be suitable for joining. For numeric columns, we calculate additional statistics: percentile values, mean, standard deviation, min and max values. We combine these features into a single vector, which we refer to as a "numerical sketch". To summarize, a numerical sketch vector consists of the following elements: [unique count, NaN count, cell width, 10th percentile, 20th percentile, …, 90th percentile, mean, standard deviation, min value, max value]. When possible, we convert date columns to timestamps and treat them as numeric columns.
@@
MinHash Sketches. For each column, we treat cell-value as a string and compute a MinHash signature from the set of cell-values.555 While it may not make sense to convert float or integer columns into MinHash signatures, we do so because it is often difficult to tell if a column is truly a float or an integer or is really a categorical value. For string columns, we also compute a MinHash signature for set of words within the column. The rationale for this second MinHash is that the words can sometimes capture its semantics (e.g., if street appears in two different columns, they maybe both about address, which is useful if one considers a union of two tables that do not necessarily share cell values). Consider the Reference Area column in the table shown in Fig. 1. The cell-values MinHash Sketch treats "Austria Vienna" as an element in the set of strings, whereas "Austria" and "Vienna" are treated as separate strings in words-based MinHash signature. Both MinHashes are concatenated into a single input vector for string columns. For numerical and date columns, only the MinHash for the cell values is included in the input vector.
The left side of Figure  1 illustrates the process of preparing the content snapshot for the table, as well as generating numerical and MinHash sketches for each individual column.
III-B Input Representation
Now that we have these sketches, which are arrays of numbers (vectors), a key question is how to adapt BERT’s architecture to consume these numeric values. Transformer models construct input embeddings from any given sentence by summing three different embeddings: 1. token embeddings—embedding of the input sentence broken into tokens, 2. token type embedding—that indicates if the token belongs to the first sentence or second, and 3. position embeddings—that indicates the position of the token in the input sentence. We adapt this to table-specific inputs in the following manner. We first generate an “input string” that is concatenation of all the table metadata and column names delimited by the separator token. The input string for table in Fig 1 is shown at the bottom right. The input embedding is constructed from this input string by summing the following different embeddings:
1. Token Embeddings. Similar to the BERT, the first embedding used for the input string is the token embeddings. We initialize the embedding weights from the pretrained BERT uncased model to leverage the extensive pretraining of BERT for its vocabulary. This is represented in Fig 1 with yellow boxes.
@@ III-B Input Representation
2. Token Position Embeddings. Considering each column analogous to a sentence in a paragraph, we re-purpose token positional embeddings to reflect a token’s position within a column name. For instance, as shown in Fig. 1 with light gray boxes, the token position of the token Area is 1 because it is the second token in that column (after Reference).
@@
3. Column Position Embeddings. Columns themselves have positions and are encoded through a new embedding layer which can range from 1 to the total number of columns, [^6] Position 0 is reserved for any tokens in the table description. as shown in green boxes in Fig. 1. The rationale for including column positions is that they sometimes do have semantics; e.g., a street address next to the city, followed by a zip code. Of course, table semantics do not change as a function of column order; nevertheless, we included a position in case it helps the model understand column semantics, with the assumption that the attentional mechanism would disregard order if necessary.
@@
4. Column Type Embeddings. The data types of columns: string, date, integer, or float are encoded through additional embedding. In order to determine column type, we made a best-case effort to parse the first 10 values of each column as dates, integers, or floats and defaulted to string if we could not convert them. In mixed-type columns, this can yield poor results, but at the very worst, at least one of the types was assigned to these columns. This is shown in Fig. 1 in orange, where string, integer, float and date types are represented with integers 1, 2, 3, and 4, resp.
5. MinHash Sketch Embeddings. This embedding encapsulates all the MinHash sketches, including those derived from cell values, words, and content snapshots. The embedding is obtained by passing the MinHash vectors through a linear transformation layer. Notably, the MinHash embedding for table meta-data tokens corresponds to the content snapshot embedding, denoted as $E_{CS}$, which is visually represented by purple boxes. For numerical columns, the MinHash embedding is the embedding of the MinHash signature derived from the set of corresponding cell values, denoted as $E_{C}$, and illustrated in red. In contrast, the MinHash embedding for string columns is constructed by concatenating the MinHash signatures from both the corresponding set of cell values and the set of words, resulting in the embedding $E_{C||W}$, shown with red and pink.
6. Numerical Sketch Embedding. Lastly, each column’s numerical sketch vector is fed into a linear transformation layer, and the output is utilized as the numerical sketch embedding.
The sum of these embeddings is used as the input embedding for the 12-layered BERT encoder model with self-attention from HuggingFace [^7] https://huggingface.co/docs/transformers/model_doc/bert. The self-attention in BERT is bi-directional: each token can attend to the tokens on both its left and the right side. This bi-directional attention helps the TabSketchFM model to disambiguate column names like Age which can mean mean person’s age in an Employee table vs age of a building in a Residential Properties table.
(a) TabSketchFM Pretraining.
(b) TabSketchFM Finetuning.
Figure 2: (a) For pretraining, the table input embedding (see Section III-B) is passed through 12-layered best encoder to generate output token embeddings. The output token embeddings are used for MLM objective with cross entropy loss( Section III-C). We initialize the weight of the BERT encoder as well as the token embeddings from the pretrained base uncased BERT model. (b) Two input tables are concatenated and passed through the pretrained TabSketchFM. The BERT pooler output (embedding of the first token) is taken and passed through a dropout and a linear layer to generate output of size N.
III-C Pretraining
@@
Dataset. Existing works generally use non-public data [18] or web tables [39] to pretrain the models. Unlike enterprise data, web tables often have few rows and columns to ease human consumption [20] and they focus on entities popular on the web (e.g., countries). Their performance may not generalize to enterprise data, where the tables have a large number of rows, the entities are often domain-specific, contain cryptic code words, and have a lot of numerical information. Therefore, we created a de-duplicated pretraining dataset of $197,254$ enterprise-like tables (CSVs) from CKAN and Socrata that are legally cleared to have the open licenses. This pretrain dataset contains $2234.5$ rows and $35.8$ columns in each table on average. About $66\%$ of columns were non-string, resembling an enterprise datalake.
@@
Data Augmentation. In the image processing literature [53], different transformations are applied over an original image to generate new image samples for training. This makes the trained model robust on different variants of the same image. We want the model to be robust to the order of columns such that shuffling columns in a table does not impact its semantics. So, we created three different versions of the table, by changing the column order, which in turn, changed the table’s content snapshot (see Section III). This increases the total number of pretraining tables to 290,948 and for each of them, we create signatures as explained in Section III.
@@
Figure 3: A combination of whole column masking, and MLM probability masking to generate up to 5 examples. The box on the left hand side is the table description and the one on the right hand side is the column names
In our scenario, for each table, we mask a single column name such that all tokens corresponding to that column are masked. This is analogous to the natural language literature’s whole word masking where all tokens corresponding to a word are masked. In addition, we used MLM probability to mask the tokens in the table description as well. As shown in Fig. 3, we generate several samples from a single table. Note that each masking gives us an example. For small tables with less than $5$ columns, we mask each of its columns one after another. However, if a table has a large number of columns, this produces a lot of examples for the same table, over-representing it. So, for large tables with more than $5$ columns, we randomly select five columns and mask them. Following this strategy over the augmented pretraining dataset, we get $730,553$ examples in training, $54,430$ examples in validation, and $58,141$ examples in test sets. We provide the experimental setup for pretraining in Section IV. Figure 2(a) shows the TabSketchFM pretraining architecture.
III-D Fine-tuning
To use our neural tabular model for identifying relevant tables in the data lakes, we finetune TabSketchFM on the LakeBench collection [17]. LakeBench is a collection of 8 benchmark that includes data from various sources such as open government data from CKAN and Socrata, economic data from the European Central Bank [54], Spider [55], and synthesized data from large knowledge graphs such as Wikidata [56]. The benchmarks cover binary classification, regression, and multi-label classification for the tasks of identifying table unionability, table joinability, and table subset. It contains 3 datasets for identifying unionable tables—TUS-SANTOS, Wiki Union, and ECB Union; 4 datasets for identifying joinable tables—Wiki Jaccard, Wiki Containment, Spider-OpenData, and ECB Join; and one dataset for identifying table subsets—CKAN Subset. Of these datasets, Wiki Jaccard, Wiki Containment and ECB Union are formulated as regression task to estimate jaccard similarity value, containment value, and the number of unionable columns, resp. (cf  [17]); ECB Join is multi-class classification to predicts the joinable columns in the table; and remaining are classification tasks. Table I summarizes the main characteristics of the datasets to show the type of tasks we used for fine tuning.
TABLE I: Cardinality of all the datasets in LakeBench, as well as search benchmarks in this paper.
@@ III-D Fine-tuning
Eurostat Subset Search 38904 2157 10.46 64.63 9.03 7.83 18.50
Wikijoin Search 46521 49.64 2.68 58.13 13.35 25.0 3.50
For each finetuning task, we create a cross-encoder [57], which evaluates if a pair of tables can be joined, unioned or are subsets of each other. Figure 2(b) shows the architecture of the TabSketchFM cross-encoder used for finetuning. For the classification and regression tasks, we use the BERT pooler output, that is the embedding of the first [CLS] token and pass it through a dropout and a linear layer to generate output of size N. For binary classification tasks, the output layer has N=2, and we use cross-entropy loss. For regression tasks, the output layer has N=1 and we use mean-squared error loss. For multi-class classification tasks, the output layer has N = number of classes and we use binary cross-entropy with logit loss for finetuning. The goal of building multiple cross encoders for the same task across different datasets was to examine whether the cross encoders would generalize across datasets; allowing re-use to new data with no additional training.
III-E TabSketchFM Overview
We propose TabSketchFM a novel sketch-based tabular representation learning architecture for dataset discovery tasks. Section III-B explains how we generate the sketches, and Section III-B describes how we adapt BERT embedding module to incorporate the numerical sketches. The model is pretrained on opendata using MLM objective to predict the columns with cross-entropy loss, described in Section III-C. After pretraining, we fine-tune the model for downstream tasks like union, join, and subset, as elaborated in Sec. III-D. Finally, we use the finetuned TabSketchFM for data discovery tasks, in the experiments. To do so, we extract the table embeddings from the finetuned TabSketchFM, and use that to create nearest neighbor indexes for search tasks. We evaluate all these aspects in the next section.
IV Experiments
Our empirical work is organized around the following research questions: